OpenStudy (anonymous):

Complete the Next Step: A partial solution to the equation 8b + 12 = 5(4 + 2b) is given below. Choose the step that is missing from the solution. 8b + 12 = 5(4 + 2b) 8b = 8 + 10b −2b = 8 b = −4 8b + 12 = 20 + 2b 8b + 12 = 20 + 10b 8b + 12 = 9 + 7b 8b + 12 = 4 + 10b

OpenStudy (anonymous):

If you start with the equation they give you, and then you look at the equation after the missing step, it is clear that they must have multiplied out that parenthesis on the right side... the parenthesis was there in the original equation, and by step 3, it is gone. So, original: 8b + 12 = 5(4 + 2b) Multiply out the right side: 8b + 12 = 20 + 10b Then you should check to see if that makes sense.... does this new form give you a way to get to step 3 which is shown as 8b = 8 + 10b ?

OpenStudy (anonymous):

so is the answer c

OpenStudy (anonymous):

I don't think so, but your answer choices aren't labeled. It looks to me, for the reasons I described above, that the missing step is: 8b + 12 = 20 + 10b That looks like your 2nd choice, so I would have thought it was "B", not "C"

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

Glad to help, hope that made sense... there's no real official process for a problem like this... just try to start solving the way they might have done and see if you can find a choice that works. If you had tried each of other options, you would have seen that none worked in that equation. Good luck...

OpenStudy (anonymous):

Solve: 2(x + 4) − (x + 1) = 10 3 1 2 0

OpenStudy (anonymous):

Similar idea here... first, multiply that parenthesis out, and also distribute the negative over the 2nd parenthesis. 2(x + 4) - (x + 1) = 10 2x + 8 - x - 1 = 10 Then collect the x terms together, and the number terms together: 2x - x + 8 - 1 = 10 And simplify: x + 7 = 10 Then finally subtract 7 from both sides to solve for x.

OpenStudy (anonymous):

so the answer is 7

OpenStudy (anonymous):

That's not what I said. x + 7 = 10 If seven plus a number is equal to 10, then what is that number?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

right :)

OpenStudy (anonymous):

sorry, I gotta take off for awhile.... good luck!